Suppose 10.00 g of an alkaline earth metal reacts with 10.0 L of water to produce 6.10 L of hydrogen gas at 1.00 atm and 25 degree celcius. Identify the metal and determine the pH of the solution.


1
Expert's answer
2020-12-25T05:11:45-0500

M + 2H2O = M(OH)2 + H2

n(H2) = p*V/R/T = (1 atm)*(6.10 L)/(0.082 L*atm/mol/K)/(25+273.15 K) = 0.2495 mol

n(M) = n(H2)

Ar(M) = m(M)/n(M) = (10 g)/(0.2495 mol) = 40.08 g/mol

The metal is calcium.

n(OH-) = 2*n(Ca(OH)2) = 2*n(Ca) = 2*0.2495 mol = 0.499 mol

c(OH-) = n(OH-)/V(solution) = (0.499 mol)/(10 L) = 0.0499 mol/L

pOH = -lg(0.0499) = 1.3

pH = 14 - pOH = 12.7